OpenStudy (anonymous):

factoring by using the difference of two squares formula 1-81z^2

OpenStudy (tkhunny):

Well, do it. \((What)^{2} = 1\) ? Likewise. \((What)^{2} = 81z^{2}\) ?

OpenStudy (anonymous):

don't know

OpenStudy (anonymous):

ok you can wtite this as (1-9z)(1+9z)

OpenStudy (anonymous):

is that fraction

OpenStudy (kc_kennylau):

@imfresh don't fool around please

OpenStudy (tkhunny):

@candyhuebner If you honestly cannot solve \((What)^{2} = 1\), then you cannot solve this problem and you are likely in the wrong class. I would recommend you go have a very honest and very clear discussion with your academic advisor.

OpenStudy (anonymous):

The formula: \[a^2-b^2 = (a-b)(a+b)\]

OpenStudy (anonymous):

\[1 - 81z^2 = a^2 - b^2\] a^2 = 1 b^2 = 81z^2 The opposite of squaring a number is taking its square root. a = sqrt(1) b = sqrt(81z^2) = sqrt(81)sqrt(z^2)